OpenStudy (anonymous):

How do I factor t - 2/t^2-3t+2)?

OpenStudy (anonymous):

I have t - 2/ ----- (t-1)(t+2)

OpenStudy (agreene):

\[\frac{t-2}{t^2-3t+2}\] is this what you're asking?

OpenStudy (anonymous):

yes.

hero (hero):

(t-2)/(t-1)(t-2) = t-1

OpenStudy (agreene):

\[\frac{1}{t-1}\] is what I'm thinking.

hero (hero):

I was just trying to show him that the t-2 cancels

OpenStudy (agreene):

Oh, i got yah know :X

OpenStudy (anonymous):

ohhh, i see.